OpenStudy (mrs.irwin):

Gaston has an annuity that pays $250 at the beginning of each quarter. If the economy grows at a rate of 4.97% annually, what is the value of the annuity if he received it in a lump sum now rather than over a period of eight years? $1,617.76 $1,698.17 $6,471.06 $6,792.67

OpenStudy (tkhunny):

Convert your interest to Quarterly. Create an appropriate quarterly discount factor. You're almost done.

OpenStudy (mrs.irwin):

how am I supposed to come up with the quarterly discount factor?

OpenStudy (mrs.irwin):

@tkhunny

OpenStudy (anonymous):

It's $6,792.67 @Mrs.Irwin

OpenStudy (mrs.irwin):

@Hannah_Waller are you 100% sure. I cannot fail?

OpenStudy (anonymous):

haha yes @Mrs.Irwin

OpenStudy (mrs.irwin):

@Hannah_Waller thank you so much!!!

OpenStudy (anonymous):

@misty1212 can u finish helping me on my problem?

OpenStudy (tkhunny):

i = 4.97% annually = 0.0497 j = i/4 = 0.012425 = Quarterly interest v = 1/(1+j) = 0.987727486 = Quarterly Discount Factor It's just a definition. Your payments now look like this, discounted to the beginning:.. 250 + 250v + 250v^2 + 250v^3 + .... Make sense?